The area of a rectangle is 240in squared. The ratio of the width to the length is 3:5. Find the Length and the Width.
Answer by scott8148(6628)

3k * 5k = 15k^2 = 240

k^2 = 16 ___ k = 4

3k = 12

5k = 20